At what temperature will a sucrose solution boil if it contains 1.55 moles of sucrose in 600mL of water? kb of water is 0.51°C/m.
1
Expert's answer
2020-10-02T14:18:52-0400

∆T=Kb×m, where m is molality (number of moles of solute per kilogramme of solvent).

If the density of water is 1g/ml, 600 ml weights 600 g or 0.6 kg.

Molality equals to

m=1.55/0.6=2.58(mol/kg)

∆T=T-T0, where T0 is a boiling point of pure water.

T=∆T+T0

T=1.32+100=101.32(°C)